[ a jar contains 55 quarters and dimes. The total value of coins is $8.50. Find the number of quarters and dimes in the jar. ]
Answer by TimothyLamb(4379)

25q + 10d = 850
q + d = 55
q = 55 - d
---
25q + 10d = 850
25(55 - d) + 10d = 850
1375 - 25d + 10d = 850
15d = 525
---
d = 35
q = 20
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php